OpenStudy (anonymous):

The biggest strawberry shortcake needed 360 square feet of strawberry topping. If this area was numerically 225 more than 3 times its length, what were the dimensions of this rectangular shortcake

OpenStudy (anonymous):

A = width * length 360 = width * length but given that 3length + 225 = 360 length = (360-225)/3 = 45 ft width = Area / length = 360 /45 = 8 ft